OpenStudy (anonymous):

Solve 7(x – 2) – 9 = 4x + 5 + 3x.

OpenStudy (anonymous):

steps are 1) remove parentheses using distributive law 2) combine like terms 3) add or subtract as needed to have variables on one side, numbers on the other 4) multiply or divide as needed to get x by itself

OpenStudy (anonymous):

7x-14-9=4x+5+3x 7x-4x-3x=14+9+5 0=28 not true

OpenStudy (aroub):

7x-14-9=4x+5+3x 7x-23=7x+5 7x-7x=23+5 0=28 no equation/solution/ true